# Out-of-Distribution Detection: A Task-Oriented Survey of Recent Advances

### v3

@@ 1. Introduction
Notably, several previous efforts have been dedicated to surveying and summarizing OOD detection in recent years.  Salehi et al. (2022) offer a comprehensive review covering Anomaly Detection, Novelty Detection, Open-Set Recognition and OOD Detection, and analyzes the relationships and distinctions among these fields.  Yang et al. (2021) propose a unified framework to discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based and reconstruction-based methods. Concurrent with our work,  Miyai et al. (2024) discuss CLIP-based methods across related domains, while our survey specifically centers on OOD detection and provides a more comprehensive coverage of recent advances in this area.  Cui and Wang (2022) conduct a survey on OOD detection from a methodological perspective but with an alternative classification criterion, including supervised, semi-supervised, and unsupervised methods. Additionally,  Lang et al. (2023); Xu and Ding (2024) review various OOD detection methods in the context of natural language processing. However, previous works focus too much on the discussion from the perspective of methods and lack an in-depth exploration from the viewpoint of task scenarios. Establishing a clear taxonomy of task scenarios can enhance a comprehensive understanding of the field and assist practitioners in selecting the appropriate method. Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a; Geng et al., 2023; Gao et al., 2023a; Du et al., 2024a; Fan et al., 2022; Wang et al., 2024a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022; Nguyen et al., 2024; Fu et al., 2024; Dai et al., 2023; Liu et al., 2024a), there is an urgent need for a comprehensive survey that incorporates the latest technologies.
@@
(§ 4.2) ,fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Model-update needed,fill=hidden-blue!30, draw=hidden-blue!80!black, [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023a), SAL  (Du et al., 2024a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Model-update free,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024a), RTL  (Fan et al., 2024), OODD  (Yang et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Large pre-trained
@@ 2. Related work
Anomaly Detection. Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009; Xing et al., 2025; Wang and Nalisnick, 2025). These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003) or system failures  (Du et al., 2017). The process involves using statistical, machine learning, or deep learning methods to model normal behavior and detect deviations. Anomaly detection is crucial in domains like cybersecurity, finance, healthcare, and manufacturing, where recognizing unusual patterns quickly can prevent significant losses or improve operational efficiency. While both AD and OOD detection aim to identify unusual or unexpected data points, compared to OOD detection, AD differs in three key aspects. First, regarding supervision, AD methods are typically unsupervised, training only on normal samples without detailed labels and detecting anomalies during testing, whereas OOD detection involves labeled category discrimination. Second, AD focuses on perceptual anomalies, where images belong to known categories but contain anomalous regions, while OOD detection targets semantic differences, with test samples from unseen categories. Additionally, although both methods commonly use AUROC (Wehler et al., 2001) for evaluation, AD specifically employs AUPRO (Bertoldo et al., 2024) to assess pixel-level detection performance.
Novelty Detection. Novelty detection (ND) focuses on identifying new or unknown data points that a model has not seen during training  (Pimentel et al., 2014). This process is essential in situations where the system needs to adapt to evolving conditions or when it’s crucial to flag new, unseen scenarios. Unlike AD, which seeks to find patterns that deviate from the norm, novelty detection aims to discover entirely new patterns. Applications include identifying new trends in social media  (Li et al., 2017), discovering new species in biological data  (Pimentel et al., 2014), or detecting new topics in document streams  (Aksoy et al., 2012). Essentially, ND deals with surprises in familiar contexts, while OOD detection deals with data from unfamiliar contexts.
Open Set Recognition. Open set recognition (OSR) extends beyond traditional classification tasks by not only classifying known classes but also recognizing when input does not belong to any of the known categories (Bendale and Boult, 2016). This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that are not present during training. Open set recognition is essential in fields like robotics  (Marsland et al., 2005), autonomous vehicles, and image recognition systems  (Amini et al., 2018), where encountering unknown objects or scenarios is common, and the system must be able to handle these gracefully. OSR is concerned with identifying new classes that are not present during training, which is generally achieved by dividing the categories of the same dataset into base classes and new classes, which means that new classes and base classes usually come from the same domain. In contrast, OOD detection focuses on identifying any data that is different from the training distribution, regardless of whether it belongs to a new class or a different domain altogether. Essentially, OSR is a subset of OOD detection, specifically aimed at classifying new class types within the same domain.
Outlier Detection. Outlier detection (OD) is the process of identifying data points that are significantly different from the majority of the data  (Singh and Upadhyaya, 2012). It is similar to AD but focuses more on the identification of individual data points rather than patterns. Outliers can arise due to variability in measurement, experimental errors, or genuinely novel variations. Applications include fraud detection, fault detection, and removing anomalous data from datasets to improve model accuracy  (Pawar et al., 2014). Compared to OOD detection, OD is more of a transductive scenario, as it inherently has access to outliers. However, classical OOD detection only encounters outliers at the time of deployment.
Zero-shot learning. Zero-shot learning is a paradigm in machine learning where the goal is to recognize objects without having seen examples of these objects during the training phase  (Romera-Paredes and Torr, 2015; Wang et al., 2019; Pourpanah et al., 2022). The fundamental challenge in zero-shot learning is to effectively transfer knowledge from seen to unseen classes  (Romera-Paredes and Torr, 2015). It has been primarily tackled by learning semantic relationships between classes through attributes or by embedding both seen and unseen classes into a shared semantic space. While both deal with unknowns during inference, zero-shot learning tries to classify new categories  (Xian et al., 2018), whereas OOD detection flags data that is anomalous or unfamiliar to the training data distribution  (Lee et al., 2018b).
Selective Classification. Selective classification, also known as reject option classification, provides a mechanism for models to abstain from making a decision when they are not sufficiently confident in their predictions  (Geifman and El-Yaniv, 2017; Chang et al., 2024). Selective classification involves a model deciding when to make a prediction based on its confidence level, effectively choosing not to predict when uncertain  (Wiener and El-Yaniv, 2011). On the other hand, OOD detection identifies data points that differ from the training distribution, aiming to flag them as unfamiliar to the model. While both methods manage uncertainty, selective classification deals with confidence in making predictions on ID data (El-Yaniv et al., 2010; Chang et al., 2024), and OOD detection focuses on recognizing and handling data that is not represented in the training set.
Misclassification (Failure) Detection. Misclassification detection aims to identify incorrect predictions on in-distribution data, exposing a model’s unreliability in its intended domain (Ahmed et al., 2022; Dong et al., 2025; Inceoglu et al., 2021; Liu et al., 2024d). This is especially important in safety-critical areas such as autonomous driving and healthcare. Failure detection generalizes this task to capture a broader range of model breakdowns, including uncertainty and data corruption. While both misclassification and OOD detection seek to improve model reliability, the former focuses on errors within the training distribution (Jaeger et al., 2023), whereas the latter targets inputs outside it.
@@ 3.1. OOD Detection Approaches with only ID Data
Reconstruction-based. The reconstruction-based approach detects OOD samples by measuring the difference between representations of data before and after reconstruction, relying on the model’s ability to accurately recover input semantics. The core idea is that OOD data contain semantic features that differ from those of ID data. By quantifying these differences, the model can effectively distinguish OOD samples. An overview is illustrated in Fig. 3 (a).
MoodCat  (Yang et al., 2022) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process. Similarly,  Zhou (2022) introduces an auxiliary module to extract activations of feature vectors, aiding the model in constraining the latent reconstruction space to filter potential OOD data. Following this, MOOD  (Li et al., 2023a) and MOODv2  (Li et al., 2024a) leverage pretraining tasks based on masked image modeling, exhibiting significant advantages in learning the internal distribution of data. PRE  (Osada et al., 2023) introduces the concept of normalized flow, combined with a penalty based on typicality to constrain reconstruction errors, which discerns differences between OOD and ID data well.
@@
Figure 3. Illustration of some subclasses within the Training-driven approaches: (a) Reconstruction-based : After extracting the feature representation $F(x)$ of the image, it is fed into a reconstruction network (e.g., VAE or DDPM) to obtain the reconstructed feature representation $F^{\prime}(x)$. By comparing the dissimilarity between the original and reconstructed images, OOD data can be identified, as OOD samples typically exhibit higher dissimilarity before and after reconstruction. (b) Prototype-based : The model learns more optimal ID prototypes during the training process to distinguish OOD data.The symbol $h$ typically denotes the features of the penultimate layer in the network. (c) OOD Synthesis : During training, real OOD data is not available; instead, OOD data is synthesized by extrapolating information from the ID data. (d) Outlier Mining : During training, real OOD data is available, but typical OOD samples are generated using this real OOD data. (e) Boundary Regularization : The model’s decision boundary is optimized by using real OOD data, typically without involving the synthesis of OOD data.
Recently, diffusion models (Ho et al., 2020) have made remarkable strides in both training stability and the quality of generated images. Leveraging diffusion models to detect OOD data has become a new research direction.  Graham et al. (2023) address this issue by introducing DDPM, utilizing it for the reconstruction of noise-disturbed images. The model employs multidimensional reconstruction error for the identification of OOD data. The information bottleneck of this model can be externally regulated. Similarly, LMD  (Liu et al., 2023a) utilizes diffusion models for OOD validation. LMD disrupts data and then employs diffusion models to reconstruct images separated from the original manifold, distinguishing OOD data by comparing differences with the original manifold. DiffGuard  (Gao et al., 2023b) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
@@
Probability-based. Probability-based approaches aim to establish probability models to describe the distribution of training data. In this domain,  Li et al. (2023b) model each ID category during training using multiple Gaussian mixture models (Rasmussen, 1999), and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.  Huang et al. (2022a) address this issue by introducing two regularization constraints. The density consistency regularization aligns the analytical density with low-dimensional class labels, and the contrastive distribution regularization helps separate the density between ID and OOD samples. Furthermore, LID (Kamkari et al., 2024) introduces a new detection criterion for the OOD detection paradox in the context of data generation by deep generative models. It measures whether data should be classified as in-distribution by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
@@
Logits-based. These methods detect OOD samples by analyzing the logits (raw outputs before softmax) from the neural network’s final layer. Logits typically represent the model’s confidence or probability for each category. LogitNorm  (Wei et al., 2022a) uses logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence. Similarly, UE-NL  (Huang and Qiao, 2023), derived from Bayesian networks, normalizes logits while simultaneously learning embeddings and uncertainty scores. It adjusts the learning intensity between samples during training, rendering the learning process robust. DML  (Zhang and Xiang, 2023) challenges the potential hindrance to OOD detection performance. Drawing on empirical insights, DML enhances OOD detection performance by decoupling logits and balancing individual components, thereby mitigating the impact of each attribute on the results. It enhances OOD detection performance by decoupling logits and scenario MaxNorm as a parameter, thereby balancing the influence of each attribute on the outcome.
@@
Lee et al. (2018a) add additional regularization terms by incorporating two supplementary terms into the original loss function. The first term compels the classifier to increase uncertainty for OOD samples, while the second term encourages the GAN to generate appropriate OOD samples. Specifically, a joint training scheme is designed to select suitable boundary samples in the low-density regions of ID samples. Similarly, VOS  (Du et al., 2022a) employs adaptive synthesis of anomaly values. It samples anomaly values from the low-likelihood regions of class-conditional distributions in the feature space, eliminating the need for reliance on external data. This enables the model to synthesize OOD data for predictions. It is worth noting that VOS supports adaptive outlier synthesis, making it easily applicable to any ID data without the need for manual intervention. Noticing the additional distributional assumptions posed by previous methods for OOD data, some researchers approach the problem from another perspective. For example, NPOS  (Tao et al., 2023) contends that, rather than modeling the feature space of outliers as a parameterized Gaussian distribution, not making any distribution assumptions about ID embedded data would endow the model with robust flexibility and generality. It selects boundary points through non-parametric density estimation based on nearest neighbors, with these boundary points situated between ID and OOD data. Additionally, SSOD  (Pei, 2024) introduces self-supervised sampling to implicitly generate OOD data. It directly samples natural OOD signals from the backgrounds of ID data images, thus overcoming limitations introduced by biases during the OOD synthesis stage.
In contrast to the implicit generation methods of OOD samples mentioned above, some approaches take a different route by directly leveraging explicit construction of OOD samples from ID samples for learning. To acquire OOD samples, CODEs  (Tang et al., 2021) initially generate seed OOD examples by slicing and stitching samples from internal distributions of different categories. These examples are then inputted into Chamfer GAN for distribution transformation, yielding high-quality OOD samples. CMG  (Wang et al., 2022a) then generates pseudo OOD data by providing class embeddings mixed as conditions to a Conditional Variational Autoencoder (CVAE), and subsequently utilizes this data to fine-tune the classifier for OOD detection. SEM  (Yang et al., 2023b) introduces a problem setting for full-spectrum OOD detection, where negative samples are generated using Mixup operations during training. It then leverages both high-level semantic information and low-level non-semantic features, such as style, to identify the source of the data. Forte (Ganguly et al., 2025) utilizes per-point statistics (such as precision, recall, and density) to capture the local geometric properties of the data manifold in the feature space and performs typicality estimation using density-based non-parametric methods (such as Gaussian mixture models). This approach helps improve the accuracy of OOD synthetic data detection. SHIFT  (Kwon et al., 2023) proposes a direct synthesis of OOD image samples based on training examples. It achieves this by employing CLIP to eliminate ID object regions within the training samples. The latent diffusion model is then utilized to substitute these regions with authentic features, all the while taking into account the contextual background. This methodology thereby establishes the model’s capability for rejection. However, ensuring the quality of generated data is often challenging. Furthermore, the quality of the generated data itself may have inherent flaws. ATOL  (Zheng et al., 2023) introduces an auxiliary task, wherein both auxiliary ID and auxiliary OOD data coexist. In a low-dimensional latent space, distinct regions are manually sought for both auxiliary ID and auxiliary OOD data, ensuring non-overlapping characteristics in the input space. Subsequently, data generation by the generator guarantees the non-overlapping properties of the input space. Finally, reliability is upheld by aligning genuine ID data with auxiliary ID data, effectively alleviating issues related to erroneously generated OOD instances. In recent studies, HamOS (Li and Zhang, 2025) employs Hamiltonian Monte Carlo (Duane et al., 1987) sampling, treating the synthesis process as sampling from a Markov chain. In the absence of OOD data, it synthesizes high-quality OOD virtual anomaly points by constructing a Markov chain (Neal, 1993). It projects the features onto a hyperspherical feature space and uses K-nearest neighbor distances to guide the generation of anomaly points. POP (Gong et al., 2025) does not require the synthesis of specific outliers, but instead directly reshapes the decision boundaries through virtual OOD prototypes, improving the model’s OOD detection capability by penalizing misclassifications.
@@
where $c\in{\{1,2,...C\}}$. Additionally, within the loss function, it enforces alignment between the embedding vectors of ID samples and class prototypes, constraining each ID class sample. This parametric OOD score can be directly obtained after training, without requiring separate estimation. CIDER  (Ming et al., 2023) builds upon SIREN by jointly optimizing two losses to enhance data discriminability, encouraging the maximization of angular distances between prototypes of different classes and the internal compactness of prototypes within the same class. The optimization process during model training pertains to the prototypes of classes. AROS (Mirzaei and Mathis, 2025) focuses more on the feature separation between ID and OOD data. It proposes using the Lyapunov stability theorem to constrain the stability of the neural network’s embedding space, ensuring that the embeddings of ID and OOD data converge to distinct stable equilibrium points. To obtain OOD data, it also employs a technique of sampling from the low-probability regions of the ID data to generate pseudo-OOD. CNC (Harun et al., 2025) proposed controlling the degree of neural collapse (NC) at different layers of the DNN network. In the Encoder, entropy regularization is introduced to mitigate NC, while in the Projector, the projection head is used to enhance the NC phenomenon, forming more compact prototype and thereby improving OOD detection performance. ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores. To address this, they propose a re-weighting optimization strategy to balance the significance of different losses. Although the ideas behind Step  (Zhou et al., 2021) are different, in the context of semi-supervised tasks, it essentially generates clusters of unlabeled ID and OOD samples through a contrastive learning process, which is conceptually similar to prototype learning.
However, PALM (Lu et al., 2024) observes that modeling each class with a single prototype may fail to capture the diversity of internal information within the data. PALM introduces a mixed-prototype strategy to optimize the modeling process, utilizing a strategy of multiple prototypes to learn informative representations for each class. By simultaneously learning class-level prototypes and contrasting inter-class prototypes, it optimizes prototype-level intra-class compactness and inter-class discriminability within the loss function. PFS (Wu et al., 2025) also introduces a clustering loss that encourages features of the same ID class to cluster more closely, thereby improving the consistency of ID data representation.
A special case: Long-tail ID data. ID data in real scenarios may present a long-tailed distribution due to the difficulty of collection and frequency of occurrence. This imbalance will strongly affect the performance of OOD detection. Many approaches are proposed to address the challenge of ID imbalance and enhance OOD detection capabilities.
OLTR  (Liu et al., 2019) addresses the robustness of tail recognition by associating visual concepts between the head and tail embeddings. It accomplishes this by dynamically adjusting embedding norms through the utilization of visual memory. The approach involves two components to enhance the robustness of long-tailed data: one utilizes direct features from standard embeddings, while the other employs memory features that store discriminative centroids derived from direct features, thereby augmenting the direct features of the images.  Mehta et al. (2022) propose a tailored complex subset mixing strategy designed specifically for handling medium-to-tail classes. This method creates mixed samples by pairing two independent samples from different tail categories and calculates the distance between the mixed sample and prototypes of specific classes. The mixing strategy is ultimately combined with prototype learning to effectively address the challenges posed by long-tail scenarios. AREO  (Sapkota and Yu, 2023) introduces a method to quantify sample uncertainty through evidential learning, akin to confidence. The entire training process is governed by an innovative multi-scheduler learning mechanism, dynamically adjusting training parameters based on the importance of different classes to ensure the model focuses on the features of minority classes, thus balancing the disparities between majority and minority classes. Overall, AREO identifies higher uncertainty in certain samples during training and dynamically adjusts parameters to mitigate overfitting to these samples.
Existing OOD detection methods predominantly assume a uniform distribution of probabilities from OOD to ID. In the presence of class imbalances,  Jiang et al. (2023a) propose an alternative strategy to enhance current OOD detection methodologies. This involves recalibrating the OOD score based on the class prior distribution of the ID and the KL divergence from the output of a pre-trained model. This approach signifies the potential to augment the robustness of OOD detection beyond the achievements of previously optimized methodologies, further fortifying the overall performance. Additionally, Open-Sampling  (Wei et al., 2022b) leverages noisy labels from the OOD dataset to rebalance the class priors of the ID training dataset. These labels are sampled from a predefined distribution that complements the original class prior distribution. COOD  (Hogeweg et al., 2024) uses the supervised model to combine individual OOD measures into a single ensemble, similar to the concept of random forests. This approach addresses the limitations of individual OOD methods and can also overcome issues related to data imbalance.

### v4

@@ 1. Introduction
Notably, several previous efforts have been dedicated to surveying and summarizing OOD detection in recent years.  Salehi et al. (2022) offer a review covering Anomaly Detection, Novelty Detection, Open-Set Recognition and OOD Detection, and analyzes the relationships and distinctions among these fields.  Yang et al. (2021) propose a unified framework to discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based and reconstruction-based methods. Concurrent with our work,  Miyai et al. (2024) discuss CLIP-based methods across related domains, while our survey specifically centers on OOD detection and provides a more comprehensive coverage of recent advances in this area.  Cui and Wang (2022) conduct a survey on OOD detection from a methodological perspective but with an alternative classification criterion, including supervised, semi-supervised, and unsupervised methods. Additionally,  Lang et al. (2023); Xu and Ding (2024) review various OOD detection methods in the context of natural language processing. However, previous works focus too much on the discussion from the perspective of methods and lack an in-depth exploration from the viewpoint of task scenarios. Establishing a clear taxonomy of task scenarios can enhance a comprehensive understanding of the field and assist practitioners in selecting the appropriate method. Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a; Geng et al., 2023; Gao et al., 2023a; Du et al., 2024a; Fan et al., 2022; Wang et al., 2024a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022; Nguyen et al., 2024; Fu et al., 2024; Dai et al., 2023; Liu et al., 2024a), there is an urgent need for a comprehensive survey that incorporates the latest technologies.
@@
(§ 4.2) ,fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Model-optimization-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023a), SAL  (Du et al., 2024a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Model-optimization-free,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024a), RTL  (Fan et al., 2024), OODD  (Yang et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Large pre-trained
@@ 2. Related work
Anomaly Detection. Anomaly detection (AD) identifies data points that deviate from normal patterns (Chandola et al., 2009), with wide applications in fraud detection (Pourhabibi et al., 2020), network security (Lazarevic et al., 2003), and system monitoring (Du et al., 2017). While both AD and OOD detection identify unusual samples, AD differs in three main aspects. First, AD is typically unsupervised, using only normal samples for training, whereas OOD detection involves labeled categories. Second, AD focuses on perceptual anomalies, while OOD detects semantic shifts from unseen categories. Lastly, although both use AUROC (Wehler et al., 2001) for evaluation, AD employs AUPRO (Bertoldo et al., 2024) for pixel-level performance. Beyond these standard metrics, practical evaluation also assesses human-centric and operational criteria, such as detection latency (Lavin and Ahmad, 2015; Sinha et al., 2024), the economic cost of errors (Xu et al., 2013; Kim et al., 2022), model interpretability (Jiang et al., 2023c; Doshi and Yilmaz, 2023), and computational footprint (Ibidunmoye et al., 2015; Huertas-Garc´ıa et al., 2023).
Novelty Detection. Novelty detection (ND) aims to identify previously unseen data points (Pimentel et al., 2014), enabling systems to adapt to new conditions or scenarios. Unlike AD, which finds deviations within known patterns, ND discovers new patterns, such as emerging social media trends (Li et al., 2017), novel species (Pimentel et al., 2014), or new document topics (Aksoy et al., 2012). In essence, ND detects surprises within familiar contexts, while OOD detection addresses data from entirely unfamiliar contexts.
Open Set Recognition. Open Set Recognition (OSR) extends traditional classification by identifying both known classes and instances from unknown categories (Bendale and Boult, 2016). Unlike OOD detection which identifies any OOD data, OSR specifically focuses on recognizing novel classes within the same domain. This makes it particularly relevant for applications like robotics (Marsland et al., 2005) and autonomous systems (Amini et al., 2018).
Outlier Detection. Outlier detection (OD) identifies individual data points that significantly deviate from the majority (Singh and Upadhyaya, 2012). Unlike OOD detection which encounters outliers only during deployment, OD is transductive with inherent access to outliers. It finds applications in fraud detection and data cleaning (Pawar et al., 2014).
Zero-shot learning. Zero-shot learning aims to recognize objects without seeing their examples during training by transferring knowledge from seen to unseen classes (Romera-Paredes and Torr, 2015; Wang et al., 2019). Unlike OOD detection which flags anomalous data (Lee et al., 2018b), zero-shot learning actively classifies new categories through semantic relationships (Xian et al., 2018).
Selective Classification. Selective classification (or reject option classification) allows models to abstain from predictions when confidence is insufficient  (Geifman and El-Yaniv, 2017; Chang et al., 2024). While OOD detection identifies samples from different distributions, selective classification focuses on managing prediction uncertainty within the training distribution  (El-Yaniv et al., 2010).
Misclassification (Failure) Detection. Misclassification detection identifies incorrect predictions on ID data (Ahmed et al., 2022; Dong et al., 2025; Inceoglu et al., 2021; Liu et al., 2024d), critical for domains like autonomous driving. It differs from OOD detection by focusing on errors within the training distribution (Jaeger et al., 2023), rather than external inputs.
@@ 3.1. OOD Detection Approaches with only ID Data
Reconstruction-based. Reconstruction-based OOD detection methods focus on measuring representational differences between original and reconstructed data. These approaches leverage the model’s capability to recover input semantics, operating on the premise that OOD samples contain distinct semantic features from ID data, as shown in Fig. 3 (a). Several notable works have advanced this direction. MoodCat  (Yang et al., 2022) employs masked image modeling with constrained synthesis based on classification outputs.  Zhou (2022) enhances detection by incorporating feature activation analysis through an auxiliary module. Building upon masked image modeling, MOOD  (Li et al., 2023a) and its successor MOODv2  (Li et al., 2024a) demonstrate improved capabilities in learning data distributions. PRE  (Osada et al., 2023) further develops this approach by combining normalized flow with typicality-based penalties to effectively distinguish between OOD and ID samples.
@@
Figure 3. Illustration of main Training-driven OOD detection approaches: (a) Reconstruction-based: Identifies OOD samples by comparing the dissimilarity between original feature representation F(x) and its reconstruction F’(x) through networks like VAE or DDPM. (b) Prototype-based: Learns optimal ID prototypes during training for OOD distinction, where h represents penultimate layer features. (c) OOD Synthesis: Generates synthetic OOD samples by extrapolating from ID data when real OOD data is unavailable. (d) Outlier Mining: Leverages available real OOD data to generate typical OOD samples during training. (e) Boundary Regularization: Optimizes model decision boundary using real OOD data without synthesis.
Recent advances in diffusion models (Ho et al., 2020) have led to improved training stability and image quality, prompting their application in OOD detection. Graham et al. (2023) introduce DDPM for reconstructing noise-corrupted images, using multidimensional reconstruction error for OOD identification and allowing external regulation of the information bottleneck. LMD (Liu et al., 2023a) also applies diffusion models, disrupting and reconstructing data to separate OOD samples from the original manifold. DiffGuard (Gao et al., 2023b) leverages pre-trained diffusion models to enhance semantic discrepancies between reconstructed OOD and original images.
@@
Probability-based. Probability-based approaches aim to establish probability models to describe the distribution of training data. In this domain,  Li et al. (2023b) model each ID category during training using multiple Gaussian mixture models (Rasmussen, 1999), and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.  Huang et al. (2022a) address this issue by introducing two regularization constraints. The density consistency regularization aligns the analytical density with low-dimensional class labels, and the contrastive distribution regularization helps separate the density between ID and OOD samples. Furthermore, LID (Kamkari et al., 2024) introduces a new detection criterion for the OOD detection paradox in the context of data generation by deep generative models. It measures whether data should be classified as ID by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
@@
Logits-based. These methods detect OOD samples by analyzing the logits (raw outputs before softmax) from the neural network’s final layer. Logits typically represent the model’s confidence or probability for each category. LogitNorm  (Wei et al., 2022a) uses logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence. UE-NL (Huang and Qiao, 2023) leverages Bayesian principles to jointly learn embeddings and uncertainty scores, normalizing logits to improve robustness in OOD detection. DML (Zhang and Xiang, 2023) further advances this area by decoupling logits and balancing key components, thereby mitigating attribute interference and enhancing detection performance.
@@
Lee et al. (2018a) and VOS (Du et al., 2022a) both synthesize OOD samples from low-density or low-likelihood regions of the ID feature space, but VOS enables adaptive outlier generation without external data or manual tuning. In contrast, NPOS (Tao et al., 2023) avoids distributional assumptions by using non-parametric density estimation to select boundary samples, enhancing flexibility and generality. SSOD (Pei, 2024) further differs by leveraging self-supervised sampling to extract OOD signals directly from ID backgrounds, addressing synthesis bias.
OOD detection methods often generate synthetic outliers from in-distribution data. Techniques include recombining features with GANs (CODEs (Tang et al., 2021)), using conditional VAEs (CMG (Wang et al., 2022a)), or applying Mixup (SEM (Yang et al., 2023b)). Alternatively, other approaches estimate data typicality by analyzing the local geometric and density statistics of the feature manifold (FORTE (Ganguly et al., 2025)). SHIFT (Kwon et al., 2023) leverages CLIP and latent diffusion models to generate OOD images by replacing ID object regions with contextually consistent features, though maintaining high sample quality remains a challenge. ATOL (Zheng et al., 2023) introduces an auxiliary task that constructs non-overlapping regions for ID and OOD data in latent space, using generated samples to reinforce this separation and aligning genuine ID data to enhance reliability. Recent studies show that HamOS (Li and Zhang, 2025) uses Hamiltonian Monte Carlo (Duane et al., 1987) to generate virtual OOD anomalies by sampling from a Markov chain (Neal, 1993) when real OOD data is unavailable. It projects features onto a hypersphere and leverages K-nearest neighbor distances to guide anomaly generation. POP (Gong et al., 2025) instead refines decision boundaries using virtual OOD prototypes and penalizes misclassifications, enhancing OOD detection without explicit outlier synthesis.
@@
where $c\in{\{1,2,...C\}}$. Additionally, within the loss function, it enforces alignment between the embedding vectors of ID samples and class prototypes, constraining each ID class sample. This parametric OOD score can be directly obtained after training, without requiring separate estimation. CIDER  (Ming et al., 2023) builds upon SIREN by jointly optimizing two losses to enhance data discriminability, encouraging the maximization of angular distances between prototypes of different classes and the internal compactness of prototypes within the same class. The optimization process during model training pertains to the prototypes of classes. AROS (Mirzaei and Mathis, 2025) enhances feature separation between ID and OOD data by applying the Lyapunov stability theorem, guiding embeddings of ID and OOD samples toward distinct stable points. Pseudo-OOD samples are generated by sampling from low-probability regions of the ID distribution. CNC (Harun et al., 2025) controls neural collapse (NC) at different network stages: entropy regularization in the encoder alleviates NC, while the projector head promotes NC, resulting in more compact prototypes and improved OOD detection. ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores. To address this, they propose a re-weighting optimization strategy to balance the significance of different losses. Although the ideas behind Step  (Zhou et al., 2021) are different, in the context of semi-supervised tasks, it essentially generates clusters of unlabeled ID and OOD samples through a contrastive learning process, which is conceptually similar to prototype learning. However, PALM (Lu et al., 2024) notes that representing each class with a single prototype often fails to capture the internal diversity of data. To address this, PALM adopts a mixed-prototype strategy, assigning multiple prototypes per class to better model informative representations. By jointly learning class-level prototypes and contrasting them across classes, PALM’s loss function encourages intra-class compactness and inter-class separability at prototype level. Similarly, PFS (Wu et al., 2025) introduces a clustering loss that further promotes tight feature clustering within each ID class, enhancing the consistency of ID representations.
A special case: Long-tail ID data. ID data in real scenarios may present a long-tailed distribution due to the difficulty of collection and frequency of occurrence. This imbalance will strongly affect the performance of OOD detection. Many approaches are proposed to address the challenge of ID imbalance and enhance OOD detection capabilities. OLTR (Liu et al., 2019), POP (Mehta et al., 2022), and AREO (Sapkota and Yu, 2023) each address long-tailed recognition from different perspectives. OLTR enhances tail class robustness by leveraging visual memory to bridge head and tail embeddings. POP focuses on medium-to-tail classes through a prototype-based mixing strategy, generating mixed samples to improve class separation. In contrast, AREO models sample uncertainty via evidential learning and dynamically adjusts training using a multi-scheduler mechanism to better balance majority and minority classes.
Most existing OOD detection methods assume a uniform probability distribution between OOD and ID samples. To address class imbalance, Jiang et al. (2023a) recalibrate OOD scores by incorporating class priors and KL divergence, improving robustness under skewed distributions. Open-Sampling (Wei et al., 2022b) rebalances ID class priors using noisy OOD labels sampled from a complementary distribution. In contrast, COOD (Hogeweg et al., 2024) ensembles multiple OOD measures via a supervised model, effectively mitigating individual method limitations and handling data imbalance.